OpenStudy (anonymous):

Bob has some 10 lb weights and some 3 lb weights. Together, all his weights add up to 50 lbs. If x represents the number of 3 lb weights and y represents the number of 10 lb weights, which equation can be used to find the number of each type of weight Bob has? A)3x – 10y = 50 B)3x = 50 – 10y C)50 + 10y = 3y D)50 + 3y = 10y

OpenStudy (anonymous):

@katie_kat01

OpenStudy (anonymous):

im thinking it's D

OpenStudy (anonymous):

well what do you think

OpenStudy (anonymous):

i think is c

OpenStudy (anonymous):

i believe you are correct but i could be wrong hold on let me look it up real quick!

OpenStudy (anonymous):

i would go with c thats what i would have told you too so i would go with c

OpenStudy (anonymous):

is there anymore i could help you with?

OpenStudy (anonymous):

holdon

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

If x is the number of 3-lb weights, how much do the 3-lb weights weigh altogether?

OpenStudy (anonymous):

Karl rides his bicycle 120 feet in 10 seconds. How many feet does he ride in A1 minute? A)2 feet B)12 feet C)720 feet D)7,200 feet

OpenStudy (anonymous):

@mathstudent55 so the first one is wrong?

OpenStudy (mathstudent55):

Yes, it's wrong. Let's do it a step at a time. How much weight is there from the 3-lb weights if you have x weights of 3 lb each?

OpenStudy (anonymous):

its D

OpenStudy (mathstudent55):

Here is a suggestion: Stop guessing. I'm trying to help you by explaining how to do this problem. If you follow my explanation, you will get it right and will understand how to do it. If you guess, you have only a 25% chance of getting it right. Once again, here is the explanation one step at a time: 1 3-lb weight weighs 3 * 1 lb 2 3-lb weights weigh 3 * 2 lb 3 3-lb weights weigh 3 * 3 lb x 3-lb weights weigh 3 * x lb That means the x number of 3-lb weights weigh 3x Ok so far?

OpenStudy (mathstudent55):

Since each 3-lb weight weighs 3-lb, the total weight of the 3-lb weights is 3 times the number of weights. Since you have x 3-lb weights, they weigh 3x lb.

OpenStudy (mathstudent55):

Now we do the same for the 10-lb weights. Each 10-lb weight weighs 10 lb. There are y of them. That means the total weight of the 10-lb weights is 10y DO you follow so far?

OpenStudy (anonymous):

omg so problems and yes thank you

OpenStudy (mathstudent55):

Ok. The 3-lb weights weigh 3x. The 10-lb weights weigh 10y. To find the total weight altogether, we add their weights and they weigh: 3x + 10y We are told the total weight is 50 lb That means 3x + 10y = 50 This is the equation we need.

OpenStudy (mathstudent55):

If you look in your choices, no equation looks exactly like this one, 3x + 10y = 50

OpenStudy (mathstudent55):

Now we look through the choices to see which equation corresponds to ours.

OpenStudy (mathstudent55):

Let's start with our equation, 3x + 10y = 50, and subtract 10y from both sides. We get: 3x = 50 - 10y Is there a choice that looks exactly like 3x = 50 - 10y?